OpenStudy (anonymous):

Can someone please tell me if I'm correct? Let f(x) be a fifth-degree polynomial, and g(x) also a fifth-degree polynomial. The degree of f(x) + g(x) must also equal 5. FALSE?

OpenStudy (anonymous):

i agree with the answer in the case where f(x) = x^5 and g(x) = -x^5+x

OpenStudy (anonymous):

Should I trust someone who calls themself a completeidiot? lol nahh but that's what I was thinking. Thank you for clearifying :D

OpenStudy (anonymous):

clarifying* and no problem

OpenStudy (anonymous):

Thanks grammer nazi